OpenStudy (dm91):

6 rows of a given length with 3 chairs left over or 8 rows of the same length with 12 more chairs

OpenStudy (dm91):

@jim_thompson5910 @KendrickLamar2014 please help

OpenStudy (anonymous):

6L + 3 = 8L + 12 Solve for L?

OpenStudy (dm91):

i think it is solve for L i am so confused on this question

OpenStudy (anonymous):

I'm telling you to solve for L.

OpenStudy (dm91):

okay

OpenStudy (anonymous):

-2l+3=12

OpenStudy (anonymous):

-2l+3=12

OpenStudy (anonymous):

-2l+3=12

OpenStudy (anonymous):

-2l+3=12

OpenStudy (anonymous):

-2l+3=12